Since helium is lighter than air, it is difficult to measure the mass of a sample, so to find out the mass of helium in a birthday balloon, a student transferred the helium to a fixed 1.50 L vessel at 25.0 °C and found the pressure to be 3.04 atm. What mass of helium is in the balloon?
Answer #1

According to ideal gas law PV = nRT p is the pressure V is the volume n is number of moles R = 0.0821 L. atm. mol.kt t is the- PV RT n = (3:04atal (1.50L) (0.0821 L-ath. mol. K) (25 +213) * n = 0-1863826 moles. mass of He= nx atomic mass of He - 0-1